OpenStudy (anonymous):

What is the solution set of |x - 4| + 7 = 4?

OpenStudy (anonymous):

x-4 = -3 x = 1 -x -4 = -3 -x = 1 x = -1 set x= +/- 1

OpenStudy (anonymous):

wrong :|

OpenStudy (anonymous):

it has no solution

OpenStudy (anonymous):

rearrange it to |x-4| = -3 this has no solution , an absolute value is never negative.